# EnScale: Temporally-consistent multivariate generative downscaling via proper scoring rules

### v2

@@ 2.3 Modeling temporal consistency
So far, we predicted each RCM day independently, ignoring the temporal context. However, our architecture can naturally be extended to generating temporally consistent samples. For time series generation, we aim to learn the distribution of the RCM data $Y_{t}$ for a given day $t$, given the GCM input $X_{t}$ on the same day and the RCM data $Y_{t-1}$ from the previous day. We restrict temporal dependence to one-day lags, similar to the Markov assumption in ML-based weather forecasting [Lam2023LearningForecasting, Price2025ProbabilisticLearning]. More long-term memory effects are potentially still present in the GCM input.
Conditioning on $Y_{t-1}$ directly is challenging because $Y_{t-1}$ is high-dimensional. Thus, we approximate the temporal dependency on the lower-dimensional scale. For this, we train a second temporal coarse model to model $Z_{t}$ given $Z_{t-1}$ and $X_{t}$. This models autocorrelation on large spatial scales, e.g. in pressure systems. During training, we use the average pooled version of the RCM data $Y_{t-1}$ as predictors for $Z_{t}$. During inference, we generate a time series with an “autoregressive rollout”: Given a sample $\hat{Z}_{t-1}$ for time step t1t-1, we predict $\hat{Z}_{t}$ with the temporal coarse model using $\hat{Z}_{t-1}$ and $X_{t}$. This is repeated iteratively for future time steps. For the initial time step $t=1$, we use one sample from the marginal coarse model, only based on $X_{1}$.
Figure 3: Time series generation with temporal consistency in EnScale-t using an autoregressive roll-out. At each time step $t$, the model predicts coarsened RCM data $\hat{Z}_{t}$ using the sample from the previous day, $\hat{Z}_{t-1}$, and the GCM data from the same day, $X_{t}$. The super-resolution model is applied to independently to each $\hat{Z}_{t}$ to generate the final high-resolution time series. For simplicity, this figure presents only one variable (precipitation), but the approach uses all variables jointly.
The resulting time series $\hat{Z}_{t}$ is then passed through the super-resolution model, without temporal dependency, to obtain a time series for $Y_{t}$ (Fig. 3). This approximation neglects temporal correlations on a local scale. We check the effect of this approximation in App. B.2, finding that the approximation does not degrade performance at all (e.g. precipitation) or by at most 2 % (e.g. surface wind).
Our approach to model temporally consistent time series approximates the distribution $p^{\iota}_{Y_{t}|X_{t},Y_{t-1}}$ with a few conditional independence statements, see App. B.2 for details.
@@ 2.4 Generative model architecture
We extend model architectures to our stochastic and high-dimensional setting by combining dense and sparse architectures. The coarse model uses a dense multilayer perceptron (MLP), following [Shen2024Engression:Regression]. However, dense layers are very memory-intensive for high-dimensional data. Thus, we design a sparse architecture for the high-dimensional super-resolution steps.
Firstly, the complex and high-dimensional super-resolution task of learning $p^{\iota}_{Y|Z}$ is split into four simpler steps, as in [Shen2025ReverseDistributions]. In particular, each step increases resolution by a factor of two in both spatial dimensions, i.e., we model $p^{\iota}_{\tilde{Z}|Z^{\prime}}$, where $Z^{\prime},\tilde{Z}$ are average-pooled versions of the target $Y$ with dimensions $2^{k}\times 2^{k}$ and $2^{k+1}\times 2^{k+1}$, respectively. The first step uses $k=3$, corresponding to the dimension of $Z$ ($8\times 8$), and $k$ increases by 1 for each step, so that the last step uses $k=6$, as $Y$ has dimension $128\times 128$.
Secondly, each conditional distribution $p^{\iota}_{\tilde{Z}|Z^{\prime}}$ is modeled with sparse local layers. The value of $\tilde{Z}$ in one location is approximated to only depend on the values in a local neighborhood grid cells in $Z^{\prime}$, with learnable, location-specific weights. This design allows the model to capture local characteristics such as topography or land-sea contrast. At the same time, by restricting interactions to local neighborhoods and by sharing non-linear transformations between locations, we keep the model size manageable even for high resolutions. Our approach combines flexibility of dense connections with efficiency of convolutional architectures.
@@
For our generative modeling task, the ML model must generate multiple outputs for one single input $X$. To achieve this, we integrate Gaussian noise into our model as a source of stochasticity. We do this separately for each subpart of our model (coarse model and each component of the super-resolution model). For the dense MLP architectures, the noise is concatenated to hidden units in each hidden layer. This facilitates learning complex representations of the random inputs and yields spatially correlated variability in the samples, in line with independently developed approaches such as [Alet2025SkillfulMarginals], [Lang2024AIFS-CRPS:Score]. For the sparse layers, the stochastic super-resolution proceeds in two stages (see Fig. 4). A first deterministic upsampling step is done separately for each target variable, and interpolates each high-resolution pixel in $\tilde{Z}$ linearly from its nearest low-resolution neighbors in $Z^{\prime}$ with learnable location-specific weights. The second step performs a stochastic refinement. The intermediate upsampled variables are stacked and concatenated with several channels of Gaussian noise in each pixel; this concatenation is used as input to another local aggregation. Each pixel again has its distinct set of linear weights, applied to inputs from a local neighborhood, using both the upsampled variables as well as the noise channels. This step enables generating spatially correlated variability and capturing multivariate dependencies. Finally, a shared MLP is used for all pixels (weight-sharing across space), where each pixel is processed independently of others, but all dimensions are considered jointly. More details on all model architectures are given in App. B.4.
@@
Figure 4: Sparse local stochastic layers from EnScale’s super-resolution model. As an example, we demonstrate modeling the distribution in an example pixel of interest (top left corner, orange); the same procedure is applied to all pixels. For each intermediate map (small arrows), light blue shaded pixels serve as inputs and orange pixels as the targets. First, a deterministic upsampling step processes each target variable (tas, pr, sfcWind, rsds) separately, linearly interpolating nearest pixels in the low-resolution input with learnable weights. Second, the intermediate upsampled variables are stacked and concatenated with noise channels. Next, each pixel again interpolates linearly from its nearest neighbors with learnable location-specific weights (this time using all variables and the noise as inputs). Finally, an MLP is applied to each pixel independently, using the same weights at each location.
Our choice of loss function ensures that the model indeed learns the correct conditional distributions, as we describe next.
@@ 2.5.1 Connection to forecasts and forecast evaluation
Our method, along with our benchmarks and evaluation techniques, is inspired by the correspondence between generative modeling and forecast evaluation. Consider the example of weather forecasts. Given the state of the atmosphere at a certain time, weather forecasts aim to predict the weather a few hours ahead. However, rather than issuing a single best estimate, they also aim to quantify uncertainties by predicting the full probability distribution of possible future weather for the next time step [Gneiting2005WeatherMethods]. Typically, this is done by running an ensemble of simulations from a numerical weather model. Afterwards, a post-processing step corrects biases and variance discrepancies of the ensemble [Vannitsem2021StatisticalWorld]. This post-processing is trained based on previous forecasts and corresponding observations. The goal is to produce a probability distribution that should be consistent with the target, i.e., the observed data.
A key challenge in evaluating such weather forecasts is that, at each time step, only a single ground truth from observations is available. Thus, tools are needed to compare a single observed value with the full predicted distribution. Averaging over multiple forecast-observation pairs then allows evaluating the quality of the distributional forecasts [Gneiting2007ProbabilisticSharpness].
@@ 2.5.2 Energy score loss
For simplicity, we first consider the unconditional generation of a target distribution $p_{Y}$ and neglect GCM predictors $X$. We denote the distribution from the ML model by $q_{Y}$. Recall that we do not have explicit access to the ML’s output distributions, but instead have samples for $p_{Y}$ and can generate samples for $q_{Y}$ via the generative ML model. We denote true data by $Y$ and two exemplary predictions from the model by $\hat{Y},\hat{Y}^{\prime}$.
@@
and equality holds if and only if $p_{Y}=q_{Y}$ [Rizzo2016EnergyDistance, SzekelyE-Statistics:Distances]. This implies that minimizing this loss ensures that the generated samples follow the target distribution.
@@
$S(q_{Y},y)\coloneqq E_{\hat{Y}\sim q_{Y}}(\lVert y-\hat{Y}\rVert)-\frac{1}{2}E_{\hat{Y},\hat{Y}^{\prime}\sim q_{Y}}(\lVert\hat{Y}-\hat{Y}^{\prime}\rVert).$
This is a multivariate generalisation of the Continuous Ranked Probability Score (CRPS) [Gneiting2007StrictlyEstimation, SzekelyE-Statistics:Distances] frequently used in numerical weather prediction. For one-dimensional $Y$, the energy score and CRPS are equivalent. The CRPS is typically applied per variable and per grid point, whereas we use the energy score jointly over all locations (and, unless stated otherwise, also across all target variables), capturing dependencies between them.
@@
2.5.3 Conditional case
We call this the EnScale-loss. We call the prediction error (first term) $\text{ES}_{\text{pred}}$ and the variability (second term) $\text{ES}_{\text{var}}$. This loss is minimized if and only if, for each predictor $x$, the predicted conditional distributions $q_{Y|X=x}$ indeed match the target conditionals $p_{Y|X=x}$. More formally, the score is minimized if and only if $q_{Y|X=x}=p_{Y|X=x}$ $p_{X}$-almost everywhere [Pacchiardi2024ProbabilisticMinimization]. As for the unconditional loss in Eq. (1), if $p_{Y|X=x}=q_{Y|X=x}$ for each $x$, it holds that $\text{ES}_{\text{pred}}=\text{ES}_{\text{var}}$. During training, we use minibatching. We describe the standard batch-wise approximation of the loss in App. B.3, where we use two samples by the generative model for each input $x$. Each batch contains data from multiple GCM-RCM pairs, such that the loss is averaged across pairs.
@@ 3.1 Data preprocessing
RCM GCM
ALADIN63 CNRM-CM5
ALADIN63 MPI-ESM-LR
CCLM4-8-17 CNRM-CM5
CCLM4-8-17 MIROC5
CCLM4-8-17 MPI-ESM-LR
REMO2015 MIROC5
RegCM4-6 CNRM-CM5
RegCM4-6 MPI-ESM-LR
@@
We use data from eight GCM-RCM pairs from the EURO-CORDEX framework [Jacob2014EURO-CORDEX:Research], see Table 1. We consider daily 2m temperature (tas), total precipitation (pr), surface wind (sfcWind) and surface downwelling shortwave (solar) radiation (rsds) from both driving GCMs and corresponding downscaled RCMs. In addition, we add sea level pressure (psl) from the GCMs as a predictor because it provides valuable information about the large scale weather pattern. The selection of climate models was based on the availability of daily surface wind data, which not all models publish. We chose a set of driving GCMs for which multiple corresponding downscaled datasets from different RCMs were available. We use daily data for the years 1971-2099, from both the historical and the RCP8.5 experiment.
@@ 3.1.3 Train-test split
The total runs span the years 1971-2099, and we use 1971-2029 and 2040-2089 as training data. A random subset of 10 % of the training data is taken as the validation set. The two test data sets are 2030-2039 and 2090-2099. The first test data set represents the “interpolation” test case, whereas the second is the “extrapolation” case with test data outside the time period covered in the training data. This yields approximately 320k data points for training and around 30k points for each test set.
@@ 3.2 Data transformation for training
First, we approximate the empirical distribution function (ECDF), $\hat{F}$, on a subsample of 10 % of the training data. The RCM’s target data from all GCM-RCM pairs are pooled, but we fit the ECDF separately for each variable and location. Applying the ECDF to the training data, again in a univariate manner, yields probabilities $\hat{F}(Y)$ that are uniformly distributed on $(0,1)$. In the case of precipitation, however, the ECDF introduces a point mass at zero rainfall and the uniformity only holds above this threshold. Second, we map to a normal distribution by computing $\Phi^{-1}(\hat{F}(Y))$, where $\Phi$ is the CDF of the standard normal distribution. This maps the probabilities $\hat{F}(Y)$ from $(0,1)$ to $\mathbb{R}$ and implies that the target values used in ML training lie in $\mathbb{R}$. This transformation ensures that all locations and variables follow the same distribution, allowing joint application of the loss to all variables.
@@ 4.1.1 Deterministic neural network
The most simple deterministic benchmark is a deterministic neural network (NN-det in the following) with an MSE loss. As input, we use five GCM variables, the four target variables plus sea level pressure. For simplicity, we have separate ML models for each target variable. More details are given in App. B.6.
@@ 4.1.2 EasyUQ
EasyUQ is a tool to convert predictions from a deterministic statistical model into probabilistic ones [Walz2024EasyOutput]. We use it to obtain predicted probability distributions from the above simple NN-det (Sec. 4.1.1). We introduce EasyUQ because, based on statistical theory, it is expected to issue good probabilistic predictions for each grid point and variable separately, as a baseline comparison for the performance of EnScale on at the level of individual locations. However, EasyUQ breaks spatial and inter-variable dependencies in the output.
EasyUQ it is applied as a post-processing step after training the NN-det, independently for each GCM-RCM pair. The NN-det’s predictions, denoted by $\hat{\mu}$, approximate the conditional mean $\mu\coloneqq E[Y|X]$. Given training data consisting of pairs of NN-det’s predictions $\hat{\mu}$ and the RCM target data $Y$, EasyUQ quantifies the uncertainty in the NN-det’s predictions. It returns full probability distributions. More details on theoretical properties and the implementation are given in App. B.6.

### v3

@@ 2.3 Modeling temporal consistency
In the previous paragraph, we considered predicting each RCM day independently, ignoring the temporal context. However, our architecture can naturally be extended to generating temporally consistent samples. For time series generation, we aim to learn the distribution of the RCM data $Y_{t}$ for a given day $t$, given the GCM input $X_{t}$ on the same day and the RCM data $Y_{t-1}$ from the previous day. We restrict temporal dependence to one-day lags. This follows the general approach in ML-based weather forecasting which typically models only low-order temporal dependence, e.g. with a two-step lag [35, 49]. More long-term memory effects are potentially still present in the GCM input.
We implement this autoregressive structure at each spatial resolution, following our stepwise modelling approach as above (Sec. 2.2). At the coarsest scale, the temporal coarse model predicts $Z_{t}$ given $Z_{t-1}$ and $X_{t}$. This models autocorrelation on large spatial scales, e.g. in pressure systems. During training, we obtain $Z_{t-1}$ via average pooling the RCM data $Y_{t-1}$. This extends naturally to further resolution steps: For each intermediate resolution $Z^{(k)}$, we predict $Z_{t}^{(k)}$ from the previous timestep at the same resolution, $Z_{t-1}^{(k)}$, and the same timestep at a coarser resolution, $Z^{(k-1)}_{t}$ (see App. C.4 for details on how the conditioning is integrated in the model architecture).
At inference time, we generate time series with an “autoregressive rollout” iteratively for each step: At the coarsest level, starting from a sample $\hat{Z}_{t-1}$ for time step t1t-1, we predict $\hat{Z}_{t}$ with the temporal coarse model using $\hat{Z}_{t-1}$ and GCM input $X_{t}$. This time series is then iteratively refined through the super-resolution steps, each predicting a higher spatial resolution, until a time series for $Y_{t}$ is generated, as shown in Fig. 3.
Figure 3: Time series generation with temporal consistency in EnScale-t using an autoregressive roll-out at each spatial resolution. At the coarsest level, the model predicts $\hat{Z}_{t}$ using the sample from the previous day, $\hat{Z}_{t-1}$, and the GCM data from the same day, $X_{t}$. This time series is then progressively refined to higher resolutions, where each step again conditions on the previous time step at the same scale and the coarser prediction from the current day.
We also tested approximating temporal consistency only on the coarse scale, i.e. using the temporal coarse model and then applying the super-resolution model independently on each day. We find that the results are very similar and this coarse approximation already models the temporal structure rather accurately in our setting.
@@ 2.4 Generative model architecture
We extend model architectures to our stochastic and high-dimensional setting by combining dense and sparse architectures. The coarse model uses a dense multilayer perceptron (MLP), following [60] (for details see App. C.4).
Secondly, we introduce a sparse architecture which models each conditional distribution $p^{\iota}_{Z^{(k+1)}|Z^{(k)}}$. Our architecture is a middle-ground between flexibility of dense connections and efficiency of convolutional architectures.
  • Breaking translation invariance: Unlike fully convolutional models, which are translation-invariant and apply the same weights everywhere, part of our model weights are specific to each location. This allows the model to capture unique local characteristics that vary by coordinate, such as topography or land-sea contrast. These local effects enter only through linear maps, which limits the risk of overfitting.
  • Structural efficiency: We keep memory size manageable even for high resolutions. First, the value of $Z^{(k+1)}$ in one location is approximated to only depend on the values in a local neighborhood grid cells in $Z^{(k)}$, with learnable, linear, and location-specific weights. This restriction reduces connections compared to dense networks. Second, we share non-linear transformations across all locations.
  • Training performance: Finally, our design reduces training times compared to deep fully convolutional networks substantially (as discussed in Sec. 5.7).
@@
For our generative modeling task, the ML model must generate multiple outputs for one single GCM input $X$. To achieve this, we integrate Gaussian noise into our model as a source of stochasticity. We do this separately for each subpart of our model (coarse model and each component of the super-resolution model). For the dense MLP architectures, the noise is concatenated to hidden units in each hidden layer. This facilitates learning complex representations of the random inputs and yields spatially correlated variability in the samples, in line with independently developed approaches such as [2], [36]. For the sparse layers, the stochastic super-resolution proceeds in two stages (see Fig. 4). A first deterministic upsampling step is done separately for each target variable, and interpolates each high-resolution pixel in $Z^{(k+1)}$ linearly from its nearest low-resolution neighbors in $Z^{(k)}$ with learnable location-specific weights. The second step performs a stochastic refinement. The intermediate upsampled variables are stacked and concatenated with several channels of Gaussian noise in each pixel; this concatenation is used as input to another local aggregation. Each pixel again has its distinct set of linear weights, applied to inputs from a local neighborhood, using both the upsampled variables as well as the noise channels. This step enables generating spatially correlated variability and capturing multivariate dependencies. Finally, a shared MLP is used for all pixels (weight-sharing across space), where each pixel is processed independently of others, but all dimensions are considered jointly. More details on all model architectures are given in App. C.4.
@@
Figure 4: Sparse local stochastic layers from EnScale’s super-resolution model, refining the low-resolution input by a factor of 2 in each dimension. As an example, we demonstrate modeling the distribution in an example pixel of interest (top left corner, orange); the same procedure is applied to all pixels. For each intermediate map (small arrows), light-blue-shaded pixels serve as inputs and orange pixels as the targets. First, a deterministic upsampling step processes each target variable (tas, pr, sfcWind, rsds) separately, linearly interpolating nearest pixels in the low-resolution input with learnable weights. Second, the intermediate upsampled variables are stacked and concatenated with noise channels. Next, each pixel again interpolates linearly from its nearest neighbors with learnable location-specific weights (this time using all variables and the noise as inputs). Finally, an MLP is applied to each pixel independently, using the same weights at each location, yielding the high-resolution output.
@@ 2.5.1 Connection to forecasts and forecast evaluation
Our method, along with our benchmarks and evaluation techniques, is inspired by the correspondence between generative modeling and forecast evaluation. Consider the example of weather forecasts. Given the state of the atmosphere at a certain time, weather forecasts aim to predict the weather a few hours ahead. However, rather than issuing a single best estimate, they also aim to quantify uncertainties by predicting the full probability distribution of possible future weather for the next time step [21]. Typically, this is done by running an ensemble of simulations from a numerical weather model. Afterwards, a post-processing step corrects biases and variance discrepancies of the ensemble [70]. This post-processing is trained based on previous forecasts and corresponding observations. The goal is to produce a probability distribution that should be consistent with the target, i.e., the observed data.
A key challenge in evaluating such weather forecasts is that, at each time step, only a single ground truth from observations is available. Thus, tools are needed to compare a single observed value with the full predicted distribution. Averaging over multiple forecast-observation pairs then allows evaluating the quality of the distributional forecasts [20].
@@ 2.5.2 Energy score loss
For simplicity, we first consider the unconditional generation of a target distribution $p_{Y}$ and neglect GCM predictors $X$. We denote the distribution from the ML model by $q_{Y}$. Recall that we do not have explicit access to the ML’s output distributions, but instead have samples for $p_{Y}$ and can generate samples for $q_{Y}$ via the generative ML model. We denote true data by $Y$ and two random predictions from the model by $\hat{Y},\hat{Y}^{\prime}$.
@@
and equality holds if and only if $p_{Y}=q_{Y}$ [52, 66]. This implies that minimizing this loss ensures that the generated samples follow the target distribution.
@@
$S(q_{Y},y)\coloneqq\underbrace{E_{\hat{Y}\sim q_{Y}}(\lVert y-\hat{Y}\rVert)}_{\text{ES}_{\text{pred}}}-\frac{1}{2}\underbrace{E_{\hat{Y},\hat{Y}^{\prime}\sim q_{Y}}(\lVert\hat{Y}-\hat{Y}^{\prime}\rVert)}_{\text{ES}_{\text{var}}}.$
This is a multivariate generalisation of the Continuous Ranked Probability Score (CRPS) [22, 67] frequently used in numerical weather prediction. For one-dimensional $Y$, the energy score and CRPS are equivalent. The CRPS is typically applied per variable and per grid point, whereas we use the energy score jointly over all locations (and, unless stated otherwise, also across all target variables), capturing dependencies between them. We call the prediction error (first term) $\text{ES}_{\text{pred}}$ and the variability (second term) $\text{ES}_{\text{var}}$.
@@
To account for the GCM inputs, our loss is adjusted for the conditional case by averaging over multiple GCM inputs $X$ and drawing samples from the generative model, which is conditioned on $x$ and the index for the GCM-RCM pair (see Sec. C.2 for more details). $\text{ES}_{\text{pred}}$ and $\text{ES}_{\text{var}}$ are defined as averages over GCM inputs $X$ accordingly.
@@ 3.1 Data preprocessing
GCM $\backslash$ RCM ALADIN63 CCLM4-8-17 REMO2015 RegCM4-6
CNRM-CM5 $\checkmark$ $\checkmark$ $\checkmark$
MIROC5 $\checkmark$ $\checkmark$
MPI-ESM-LR $\checkmark$ $\checkmark$ $\checkmark$
For training and evaluation, we use data from eight GCM-RCM pairs from the EURO-CORDEX framework [31], see Table 1. We consider daily 2m temperature (tas), total precipitation (pr), surface wind (sfcWind) and surface downwelling shortwave (solar) radiation (rsds) from both driving GCMs and corresponding downscaled RCMs. In addition, we add sea level pressure (psl) from the GCMs as a predictor because it provides valuable information about the large scale weather pattern. The selection of climate models was based on the availability of daily surface wind data, which not all models publish. We chose a set of driving GCMs for which multiple corresponding downscaled datasets from different RCMs were available. We use daily data for the years 1971-2099, from both the historical and the RCP8.5 experiment.
We also tested extrapolation to GCMs unseen during training: For this, we use ALADIN63 – HadGEM2-ES and CCLM4-8-17 – CanESM2 in addition to the eight training pairs (see Sec. 7).
@@ 3.1.3 Train-test split
The total runs span the years 1971-2099, and we use 1971-2029 and 2040-2089 as training data. We pool the training data across all GCM-RCM pairs and then select a random subset of 10 % of this data as the validation set. The two test data sets are 2030-2039 and 2090-2099. The first test data set represents the “interpolation” test case, whereas the second is the “extrapolation” case with test data outside the time period covered in the training data. This yields approximately 320k data points for training and around 30k points for each test set.
@@ 3.2 Data transformation for training
First, we approximate the empirical distribution function (ECDF), $\hat{F}$, on a subsample of 10 % of the training data. The RCM’s target data from all GCM-RCM pairs are pooled, but we fit the ECDF separately for each variable and location. Applying the ECDF to the training data, again in a univariate manner, yields probabilities $\hat{F}(Y)$ that are uniformly distributed on $(0,1)$. In the case of precipitation, however, the ECDF introduces a point mass at zero rainfall and the uniformity only holds above this threshold. Second, we map to a normal distribution by computing $\Phi^{-1}(\hat{F}(Y))$, where $\Phi$ is the CDF of the standard normal distribution. This maps the probabilities $\hat{F}(Y)$ from $(0,1)$ to $\mathbb{R}$. This transformation ensures that all locations and variables follow the same distribution, allowing joint application of the loss to all variables.
@@ 4.1.1 Deterministic neural network
The most simple deterministic benchmark is a deterministic neural network (NN-det in the following) with an MSE loss. As input, we use five GCM variables, the four target variables plus sea level pressure. For simplicity, we have separate ML models for each target variable. The model architecture is the same as EnScale’s coarse model, using a dense MLP but without additional noise input. More details are given in App. C.6.
@@ 4.1.2 EasyUQ
EasyUQ is a tool to convert predictions from a deterministic statistical model into probabilistic ones [72]. We use it to obtain predicted probability distributions from the above simple NN-det (Sec. 4.1.1). We introduce EasyUQ because, based on statistical theory, it is expected to issue good probabilistic predictions for each grid point and variable separately, as a baseline comparison for the performance of EnScale on at the level of individual locations. However, EasyUQ breaks spatial and inter-variable dependencies in the output.
EasyUQ it is applied as a post-processing step after training the NN-det, independently for each GCM-RCM pair. The NN-det’s predictions, denoted by $\hat{\mu}$, approximate the conditional mean $\mu\coloneqq E[Y|X]$. Given training data consisting of pairs of NN-det’s predictions $\hat{\mu}$ and the RCM target data $Y$, EasyUQ quantifies the uncertainty in the NN-det’s predictions. It returns full probability distributions. More details on theoretical properties and the implementation are given in App. C.6.